When antimatter is mixed with matter, they annihilate to energy according to E = mc2. Assuming 3.5 kg of antimatter mix with matter. What’s the approximate total amount of energy is released in kilotons of TNT? (Note: 20 kiloton TNT = 8.4 × 1013 J).

.A 70.0 megatons

B 0.3 kilotons

C 140.0 megatons

D 1.4 megatons

E 28.0 kilotons
Answer #1

Whenm= 3.5 kg antimatter is mixed with matter

The energy released = mc^2 where c=speed of light =3*10^8 m/S

So the amount of energy released= 3.5× (3*10^8 )^2J = 3.15×10^17 J

So amount of TNT = {(3.15×10^17)/(8.4×10^13) }×20 kiloton = 75×10^6 tons = 75 megatons ~ 70megatons

Answer :A. 70 megatons